OpenStudy (anonymous):

Find the Integral of {e^x(1 + xInx)}/x, for x>1

OpenStudy (anonymous):

it doesn't have a simple form. integral ( {e^x(1 + xInx)}/x dx) = integral ( {e^x}/x dx ) +integral( {lnx} dx ) = integral ({e^x}/x dx) + (x-1)* lnx unfortunately integral ({e^x}/x dx) cannot be broken down into elementary functions, and is called the exponential integral, as a function: Ei(x) so if you permit the exponential integral as an elementary function then: Integral ( {e^x(1 + xInx)}/x dx) = Ei(x) + (x-1)* lnx +c